OpenStudy (anonymous):

Using a directrix of y = -3 and a focus of (2, 1), what quadratic function is created?

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@this_emo_chic1023 @triciaal

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

@One098

OpenStudy (one098):

@TuringTest

OpenStudy (anonymous):

@iambatman @TuringTest @Preetha @iGreen

OpenStudy (igreen):

Jk :p

OpenStudy (igreen):

We have the formula: \((x-a)^2 + b^2 - c^2 = 2(b - c)y\) Where (a, b) is the focus, and 'c' is the directrix. Here, the focus is (2, 1) and the directrix is -3. So can you plug in a = 2, b = 1, and c = -3? @Brostep0s

OpenStudy (anonymous):

\[(x-2)^2 + 1^2 +3^2 = 2(1+3)y\]

OpenStudy (igreen):

Yep, you got it!

OpenStudy (igreen):

I guess we can simplify this.. Let's simplify \((x-2)^2\) first. \((x - 2)^2 \rightarrow x^2 - 2^2 \rightarrow x^2 + 4\) Plug that back in: \(x^2 + 4 + 1^2 + 3^2 = 2(1 + 3)y\) Now let's simplify \(2(1 + 3)y\) \(2( 1 + 3) y \rightarrow [(2 \times 1) + (2 \times 3)]y \rightarrow (2 + 6)y \rightarrow 8y\) Plug that back in: \(x^2 + 4 + 1^2 + 3^2 = 8y\) Can you simplify the exponents and combine the like terms?

OpenStudy (igreen):

Um..what are your options..? Because I'm not sure which form they want it in..but I'm pretty sure it's standard form(\(ax + bx^2 + c\)).

OpenStudy (anonymous):

\[ f(x) = \frac{ 1 }{ 8 }(x - 2)2 - 1\]

OpenStudy (anonymous):

That's how they want it/

OpenStudy (igreen):

Oh..that's vertex form..

OpenStudy (igreen):

Okay, well can you simplify \(x^2 + 4 + 1^2 + 3^2 = 8y\)?

OpenStudy (anonymous):

y=\[y=\frac{ x^2 }{ 8 } +\frac{ 7 }{ 4 }\]

OpenStudy (igreen):

................

OpenStudy (anonymous):

It's wrong, isn't it?

OpenStudy (igreen):

\(1^2 \rightarrow 1 \times 1 \rightarrow 1\) \(3^2 \rightarrow 3 \times 3 \rightarrow 9\) Plug those back in: \(x^2 + 4 + 1 + 9 = 8y\) Now can you add 4 + 1 + 9 to simplify it completely?

OpenStudy (anonymous):

14?

OpenStudy (igreen):

Yep..so we have: \(x^2 + 14 = 8y\) What you did was correct though..you divided 8 to both sides..giving us: \(y = \dfrac{x^2}{8} + \dfrac{14}{8}\) And \(\dfrac{14}{8}\) simplifies to \(\dfrac{7}{4}\)..which gives us: \(y = \dfrac{x^2}{8} + \dfrac{7}{4}\) To write this in function notation, we just switch \(y\) with \(f(x)\): \(f(x) = \dfrac{x^2}{8} + \dfrac{7}{4}\)

OpenStudy (anonymous):

THAT'S EXACTLY WHAT I PUT THE FIRST TIME!

OpenStudy (igreen):

I know that..I said you were correct..lol.

OpenStudy (anonymous):

oh. Wow... If we were doing this in public, we would be having one of those awkward silences in which a homosexual baby was born. Congrats.

OpenStudy (igreen):

........................

OpenStudy (igreen):

Erm...well I think I'm doing something wrong..hold on.

OpenStudy (anonymous):

NEVERMIND! You broke the silence lol

OpenStudy (anonymous):

This is my last question for the 2nd part of my segment exam, so if we could move this along, that would be great lol

OpenStudy (igreen):

Okay, well we can write: \(\sqrt{(x + 2)^2 + (y + 1)^2} = \sqrt{(y - 3)^2}\) ^^^^^^^^^^^^^^^^^^ This is the distance formula. Where '2' is the x-value of our focus, and '1' is the y-value of our focus, and '-3' is our directrix. Then we write: \((x + 2)^2 + y^2 + 4y + 4 = y^2 - 6y + 9\) The \(y^2\)'s cancel out: \((x + 2)^2 + 4y + 4 = - 6y + 9\) Subtract 4y to both sides: \((x + 2^2) + 4 = -10y + 9\) Subtract 4 to both sides: \((x + 2)^2 = -10y + 5\) Now we factor \(-10y + 5\) which gives us \(5(-2x + 1)\) So we have: \((x + 2)^2 = 5(-2x + 1)\) I don't know what to do from here ;-;

OpenStudy (igreen):

@ganeshie8 @TuringTest

OpenStudy (anonymous):

Idk, I'm just gonna guess. Thanks for the help

OpenStudy (igreen):

Wait..what are your options?

OpenStudy (anonymous):

Too late, & I got it wrong, but I got a B+ on the exam.

OpenStudy (igreen):

Oh.. B+ is very good..congratz.

OpenStudy (igreen):

@cwrw238 @zepdrix @iambatman